OpenStudy (anonymous):

Blue jeans, Inc. sells jeans that cost $15.99 for a selling price of $42.95. What is the percent of markup based on cost? (Round to nearest hundredth percent.)

OpenStudy (tkhunny):

You do have to be careful with this question. There are two popular ways to define markup. You must be SURE you know which one you want. Markup: 42.95 - 15.99 = 26.96 #1 % of the Wholesale Price. Q: How does my profit compare to my cost? % Markup: 26.96 / 15.99 = 168.61% <== A value so calculated can be very large. #2 % of the Retail Price. Q: How much of my revenue is markup? % Markup: 26.96 / 42.95 = 62.77% <== A value so calculated MUST be less than unity (1.00). Did the problem statement give you enough information to choose the right one?